OpenStudy (anonymous):

Hooke's Law. The distance d when a spring is stretched by a hanging object varies directly as the weight w of the object. If the distance is 23 cm when the weight is 3 kg, what is the distance when the weight is 5 kg? I need the answer badly and I do not understand the problem.

OpenStudy (ashwinram):

distance d is directly proportional to weight w so, d = k * w (where k is some constant you don't know yet)

OpenStudy (ashwinram):

when d = 23, w = 3 so, 23 = k * 3 which means k = 23/3

OpenStudy (ashwinram):

now, you have to find d when w = 5 d = k * w = 23/3 * 5

OpenStudy (ashwinram):

another way of thinking about it: 3 kg stretches the spring by 23 cm so 5 kg will stretch it even more

OpenStudy (ashwinram):

3 kg stretches it by 23 cm in other words, each kg stretches it by 23/3 cm so, 5 kg will stretch it 5 times that much = 5 * 23/3

OpenStudy (anonymous):

ok thanks